### v1

# HOMURA: Taming the Sand-Glass for Time-Constrained LLM Translation via Reinforcement Learning

Abstract
Large Language Models (LLMs) have achieved remarkable strides in multilingual translation but are hindered by a systemic cross-lingual verbosity bias, rendering them unsuitable for strict time-constrained tasks like subtitling and dubbing. Current prompt-engineering approaches struggle to resolve this conflict between semantic fidelity and rigid temporal feasibility. To bridge this gap, we first introduce Sand-Glass, a benchmark specifically designed to evaluate translation under syllable-level duration constraints. Furthermore, we propose HOMURA, a reinforcement learning framework that explicitly optimizes the trade-off between semantic preservation and temporal compliance. By employing a KL-regularized objective with a novel dynamic syllable-ratio reward, HOMURA effectively "tames" the output length. Experimental results demonstrate that our method significantly outperforms strong LLM baselines, achieving precise length control that respects linguistic density hierarchies without compromising semantic adequacy.
1 Introduction
With the rapid advancement of Large Language Models (LLMs), their multilingual capabilities have achieved unprecedented semantic adequacy and fluency in Neural Machine Translation (NMT) Hendy et al. (2023); Jiao et al. (2023); Zhu et al. (2024). However, this linguistic prowess is accompanied by a systemic cross-lingual verbosity bias Briakou et al. (2024); Manakhimova et al. (2024). As illustrated in Figure 1, LLM-generated translations tend to be significantly longer than their source utterances, a "chattiness" that persists even when semantic content remains constant. While acceptable in general text generation, this expansion becomes a critical failure mode in time-constrained scenarios such as subtitles, video dubbing, and simultaneous interpretation,where translation length constitutes a rigid temporal budget that strictly limits the output duration Karakanta et al. (2020b); Jørgensen and Mengshoel (2025); Gaido et al. (2024). To operationalize the temporal budget, we approximate the available speaking time with a cross-lingually comparable syllable budget.
Figure 1: Illustration of cross-lingual verbosity bias. Unlike forward metrics, the Roundtrip Expansion Ratio $\rho_{\text{rtp}}$ isolates model-induced redundancy from linguistic density shifts. A value of $\rho_{\text{rtp}}>1.0$ indicates systemic inflation despite constant semantic content. Formal definition and diagnostic study are provided in Section 2.
Viewed through an information-theoretic lens Shannon (1948); Davisson (1972), these temporal budgets act as a strict rate constraint: with a limited number of target syllables available, the model must pack maximal meaning into minimal length, implicitly operating near a rate-distortion limit on the achievable fidelity compression trade-off. In our setting, the rate is the syllable budget, and distortion can be understood as the loss in meaning or adequacy relative. Unfortunately, off-the-shelf LLMs struggle to respect these limits. This failure is deeply rooted in the post-training alignment process, where learned reward models frequently exhibit a "length bias" that conflates verbosity with quality Saito et al. (2023); Singhal et al. (2023). Furthermore, cross-lingual typological differences systematically skew length ratios, transforming length control from a superficial formatting task into a first-class modeling challenge.  [^1] Detailed related works can be found in appendix A.
To address this fundamental mismatch between LLM capabilities and temporal constraints, we argue that length control should be treated as an optimization objective rather than a post-hoc adjustment. To this end, we first operationalize the problem by constructing Sand-Glass, a dedicated benchmark for time-constrained media localization. The name reflects the rigid temporal budget that "drains" as the model produces redundant tokens. Unlike generic translation datasets, Sand-Glass incorporates syllable-based duration proxies calibrated by Information Density (ID), derived from real-world colloquial subtitles enabling precise evaluation of length compliance.
Building on this testbed, we propose HOMURA (Hard Optimization for Multilingual Utterance Reduction and Alignment), a reinforcement learning framework designed to “tame the sandglass” by explicitly regulating translation length without requiring supervised compression datasets. Instead of relying on brittle prompts, HOMURA optimizes the quality compression trade-off directly. We design a novel reference free reward function that combines a dynamic syllable ratio penalty with a semantic fidelity reward, encouraging the model to condense information density without sacrificing core meaning.
Our core contributions are as follows:
  • Diagnostic Analysis: We systematically quantify the verbosity bias in LLM-based translation under temporal constraints, providing cross-lingual diagnostics (e.g., the $\rho_{\text{rtp}}$ metric) that reveal the limitations of unconstrained models as visualized in Figure 1.
  • Benchmark Construction: We introduce Sand-Glass, a specialized dataset tailored for time-constrained translation, featuring syllable-level duration budgets to simulate real-world dubbing and subtitling scenarios.
  • Methodological Innovation: We propose HOMURA, a KL-regularized RL framework that dynamically balances strict length constraints with semantic preservation, significantly outperforming strong LLM baselines in both compliance and quality.
2 Pilot Study: Benchmark and Verbosity Bias
Figure 2: Forward vs. backward expansion relationship (colored by round-trip ratio $\rho_{\text{rtp}}$). The red curve represents the unity baseline $\rho_{\text{rtp}}=1.0$ ($y=1/x$). Points above this curve indicate systemic model-induced length inflation. Unconstrained models show a median $\rho_{\text{rtp}}>1.10$.
2.1 The Sand-Glass Benchmark
To diagnose cross-lingual verbosity bias under temporal constraints, we introduce Sand-Glass, a benchmark derived from colloquial video transcripts. Unlike generic datasets, it incorporates syllable-based duration proxies to strictly test time-sensitive compliance. The construction involves duration-aware segmentation based on speech pauses and core event extraction for semantic evaluation. We refer readers to Appendix D for detailed protocols on dataset filtering and event extraction.
2.2 Syllable as Duration Proxy
Our use of syllables as duration proxies is grounded in the Iso-Information Principle. Pellegrino Pellegrino et al. (2011) demonstrated that while Syllable Rate (SR) vary drastically across languages (e.g., Spanish $7.82\sigma/s$ [syllables/s] vs. Mandarin $5.18\sigma/s$), the universal Information Rate remains nearly constant at $\approx 39$ bits/s Coupé et al. (2019).Recent findings further corroborate this by quantifying a universal surprisal–duration trade-off (Pimentel et al., 2021), showing that languages systematically adjust segment duration to smooth information transmission.
Language Syllable Rate Info. Density Expansion
($\sigma/s$) (Normalized) (from Zh)
Mandarin (Zh) $5.18\pm 0.15$ $0.94$ $1.00\times$
English (En) $6.19\pm 0.16$ $0.91$ $\approx 1.03\times$
German (De) $5.97\pm 0.19$ $0.79$ $\approx 1.19\times$
Spanish (Es) $7.82\pm 0.16$ $0.63$ $\approx 1.49\times$
Table 1: Cross-linguistic statistics adapted from Pellegrino Pellegrino et al. (2011). Lower density necessitates higher expansion factors when translating from Mandarin.
As shown in Table 1, this trade-off implies that low-density languages must utilize higher syllable counts to convey equivalent content. Translating from a high-density source (Mandarin, ID=0.94) to a lower-density target (e.g., Spanish, ID=0.63) necessitates a linguistically required expansion. A direct 1:1 mapping would result in $\approx 33\%$ information loss (derived from the density ratio $\frac{0.63}{0.94}$). Thus, our constraints are non-uniform and calibrated to these density ratios to distinguish legitimate expansion from model verbosity.
2.3 Quantifying Cross-Lingual Verbosity Bias
Standard metrics like the Forward Expansion Ratio defined as $\rho_{fwd}=\sigma(y)/\sigma(x)$ for a source $x$ and its translation $y$ often confound linguistic necessity with model hallucinations. For instance, a high $\rho_{fwd}$ in Spanish might simply reflect its naturally lower information density rather than model-induced redundancy.
To isolate systemic bias, we introduce the Roundtrip Expansion Ratio ($\rho_{rtp}$). By translating the target output $y$ back into the source language ($\hat{x}$) and comparing its syllable count to the original input $x$, we define:
$\rho_{rtp}=\rho_{fwd}\cdot\rho_{bwd}=\frac{\sigma(y)}{\sigma(x)}\cdot\frac{\sigma(\hat{x})}{\sigma(y)}=\frac{\sigma(\hat{x})}{\sigma(x)},$ (1)
where $\sigma(\cdot)$ denotes the syllable count of a sequence. Theoretically, since this comparison is performed within the same source language, $\rho_{rtp}$ should gravitate towards 1.0. Any value where $\rho_{rtp}>1.0$ serves as a density-invariant indicator of model-induced inflation.
Empirical analysis on Sand-Glass (Figure 2) confirms that this verbosity is systemic rather than anecdotal. By plotting the joint distribution of expansion ratios, we observe a pervasive upward shift from the unity baseline ($\rho_{rtp}=1.0$), with the vast majority of segments clustered in the “inflation zone.” This concentration demonstrates a consistent model-induced redundancy that persists across various frontier models and language pairs (see Appendix L for statistical breakdowns).
Calibrated Target Bounds.
To bridge this gap, we define Target Syllable Bounds ($\mathcal{B}_{L}$) by contrasting model inflation ($\mu_{LLM}$) against Theoretical Expansion ($\mu_{theo}$). As shown in Table 2, LLMs produce significant surplus bias ($\Delta_{bias}$). We establish $\mathcal{B}_{L}$ to enforce a strict condensation regime. Notably, we lower these targets beyond theoretical baselines to address real-world dubbing and subtitling complexities, specifically to buffer against rapid speech and synchronization constraints. Accordingly, we configure the bounds as follows: English ($\mathcal{B}_{En}\in[0.8,0.9]$) is set below parity to force active summarization; German ($\mathcal{B}_{De}\in[0.9,1.0]$) maintains relative pressure despite lower density; and Spanish ($\mathcal{B}_{Es}\in[1.0,1.1]$) significantly reduces the unconstrained baseline of $1.84\times$ to ensure intelligibility under strict time limits while respecting its information density.
Target Theoretical LLM Baseline Bias Ours (Target)
Lang ($\mu_{theo}$) ($\mu_{LLM}$) ($\Delta$) ($\mathcal{B}_{L}$)
En $1.03\times$ $1.35\times$ +0.32+0.32 $\mathbf{0.8}\sim\mathbf{0.9}$
De $1.19\times$ $1.59\times$ +0.40+0.40 $\mathbf{0.9}\sim\mathbf{1.0}$
Es $1.49\times$ $1.84\times$ +0.35+0.35 $\mathbf{1.0}\sim\mathbf{1.1}$
Table 2: Comparison of expansion ratios. Target Bounds ($\mathcal{B}_{L}$) are set to remove model bias ($\Delta$) and enforce semantic condensation.
3 Method
We formulate syllable-level control as a standard controllable text generation taskLi et al. (2024); Gu et al. (2025), where our objective is to maximize translation quality while satisfying the prescribed syllable-ratio constraint. Our method is illustrated in Figure 3.
Figure 3: Overview of our HOMURA framework. Given a source utterance $x$ and a syllable-ratio budget $c$, we optimize a KL-regularized policy $\pi_{\theta}(y\!\mid\!x,c)$ with GRPO. The reward combines a length-compliance term $R_{\mathrm{len}}$ and a quality term $R_{\mathrm{qual}}$ (rubric-based or reason-based), encouraging concise yet faithful translations.
3.1 KL-Regularized Objective for Controllable Generation
We study conditional generation with an input $x$ and a control condition $c$. Our objective is to learn a policy $\pi_{\theta}(y\mid x,c)$ that satisfies $c$.
Target Distribution via a Controlled Posterior.
A natural target distribution for controllable generation is the controlled posterior
$P(y\mid x,c)\propto P(c\mid x,y)\,P_{0}(y\mid x),$ (2)
where $P(c\mid x,y)$ measures how well an output $y$ meets the control condition, and $P_{0}(y\mid x)$ denotes a fixed reference model that captures the unconstrained translation distribution given the input $x$.
KL Projection.
We fit $\pi_{\theta}(\cdot\mid x,c)$ by projecting it onto the target distribution:
$\pi_{\theta}^{\star}(\cdot\mid x,c)=\arg\min_{\pi_{\theta}}\ \mathrm{KL}\!\left(\pi_{\theta}(\cdot\mid x,c)\ \|\ P(\cdot\mid x,c)\right).$ (3)
Ignoring constants independent of $\theta$ and using Eq. (2), this objective is equivalent to maximizing
$J(\theta)$ $=\mathbb{E}_{y\sim\pi_{\theta}}\Big[\log P(c\mid x,y)-\beta\log\tfrac{\pi_{\theta}(y\mid x,c)}{P_{0}(y\mid x)}\Big],$ (4)
where $\beta>0$ is a temperature (or weighting) coefficient that controls the strength of KL regularization. In practice, we model $\log P(c\mid x,y)$ with a task reward $R(x,c,y)$.
Eq. (4) can be viewed as a KL-regularized reinforcement learning objective: $R(x,c,y)$ encourages satisfying the control condition, while the KL term discourages drifting away from the base model $P_{0}$. The coefficient $\beta$ controls the trade-off between control strength and faithfulness.
Optimization with GRPO.
We optimize Eq. (4) using GRPO (Shao et al., 2024), a lightweight PPO-style algorithm for sequence-level rewards. For each $(x,c)$, we sample multiple candidates from the previous policy $\pi_{\theta_{\text{old}}}$, compute rewards and group-normalized advantages, and update $\pi_{\theta}$ with a clipped policy objective together with the KL regularization toward $P_{0}$.
3.2 Reward Design
In our setting, the control condition $c$ corresponds to satisfying a strict length budget while maintaining semantic consistency with the source. This requires a reward that simultaneously encourages length compliance and discourages meaning drift. We interpret $\log P(c\mid x,y)$ in Eq. (4) as a task reward, denoted by $R(x,c,y)$, and instantiate it as:
$R(x,c,y)=\lambda_{\mathrm{len}}\,R_{\mathrm{len}}(x,y)+\lambda_{\mathrm{qual}}\,R_{\mathrm{qual}}(x,y),$ (5)
where $x$ is the source utterance, $y$ is the generated translation, and $\lambda_{\mathrm{len}},\lambda_{\mathrm{qual}}\geq 0$ control the trade-off between temporal alignment and semantic fidelity.
3.2.1 Dynamic Syllable-ratio Reward
To ensure the translation length aligns naturally with the source speech across varying scales, we propose a dynamic syllable-ratio reward. The syllable ratio $\rho$ for a source-target pair $(x,y)$ is defined as:
$\rho(x,y)=\frac{\sigma(y)}{\max(\sigma(x),1)},$ (6)
where $\sigma(\cdot)$ is the syllable count of a sequence.
Recognizing that strict length constraints must be sensitive to both linguistic density and utterance length, we implement a dynamic acceptance interval $[L(x),R(x)]$. We initialize the baseline bounds $L_{0},R_{0}]$ using the language-specific Target Bounds $\mathcal{B}_{L}$ set in Section 2.3. This ensures the optimization target respects the cross-lingual density hierarchy established in our theoretical analysis.
However, for shorter utterances, the “granularity” of syllable counting introduces high variance (discretization noise), making rigid bounds brittle. To address this, for utterances where the source length $\sigma(x)$ is below the corpus mean $\mu$, we define a scaling factor $\gamma(x)$ to adaptively relax the lower threshold:
$\gamma(x)=\alpha_{1}+\alpha_{2}\left(\frac{\sigma(x)}{\mu}\right)^{1/2},$ (7)
where $\alpha_{1},\alpha_{2}$ are hyperparameters governing the degree of relaxation. The final dynamic bounds are formulated as $L(x),R(x)]=[L_{0}\cdot\gamma(x),\,R_{0}]$.
To facilitate stable policy optimization, we avoid sparse or binary signals in favor of a smoother reward landscape. The reward function $R_{\mathrm{len}}(x,y)$ evaluates compliance with these dynamic boundaries via a squared exponential decay:
$R_{\mathrm{\rho}}=\begin{cases}1&\text{if }\rho\in[L(x),R(x)]\\ \exp(-k\delta^{2})&\text{otherwise}\end{cases}$ (8)
where $\delta(x,y)=\max\bigl(|\rho(x,y)-L(x)|,|\rho(x,y)-R(x)|\bigr)$ denotes the deviation. This non-linear formulation provides a soft penalty that effectively guides the model toward the density-calibrated region while preventing training instability associated with hard-truncation rewards. In Appendix F, we further investigate the effectiveness of a static syllable-ratio reward.
3.2.2 Rubric-based Quality Reward
We adopt a two-part quality rubricHashemi et al. (2024); Liu et al. (2023), covering (i) semantic fidelity and (ii) linguistic well-formedness, and implement it as the following reward components.
Semantic Fidelity Reward
Inspired by the probabilistic duality in (He et al., 2016), we enforce content preservation via a reconstruction-based reward. We employ a frozen DeepSeek-V3 (DeepSeek-AI, 2024) to back-translate the hypothesis $y$ into $\hat{x}=f_{bt}(y)$. To measure semantic retention, we compute the cosine similarity S(x,x^)S($\hat{x}$) between embeddings of the source $x$ and reconstruction $\hat{x}$ extracted by Qwen-Embedding-0.6B (Zhang et al., 2025):
$S(x,\hat{x})=\frac{\mathbf{v}_{x}\cdot\mathbf{v}_{\hat{x}}}{\|\mathbf{v}_{x}\|\|\mathbf{v}_{\hat{x}}\|}$ (9)
The final reward is clipped to stabilize optimization:
$R_{bt}(x,y)=\text{clip}(S(x,\hat{x}),\tau_{min},\tau_{max})$ (10)
Linguistic Well-formedness Reward
To prevent the telegraphic outputs often associated with aggressive compression, we explicitly penalize linguistic degradation. We utilize a frozen DeepSeek-V3 as a judge to evaluate translations against a rubric covering: (i) grammaticality, (ii) readability, and (iii) language consistency (e.g., no code-switching). The model outputs a binary decision $s_{flu}\in\{0,1\}$, serving directly as the reward:
$R_{flu}(x,y)=s_{flu}$ (11)
Full prompting details are in Appendix I.
3.2.3 Reason-based Quality Reward
Beyond the simple compositional rewards, we explore an end-to-end generative reward model (GenRM)Mahan et al. (2024); Liang et al. (2025) that evaluates translation quality via explicit reasoning. GenRM is trained using structured Chain-of-Thought (CoT) annotations synthesized by Gemini-2.5-Pro (Comanici et al., 2025). To ensure high-quality reasoning, we filter the raw CoT data using the aforementioned rubric-based method. The model is instantiated from Qwen3-32B (Yang et al., 2025) and optimized via supervised fine-tuning (SFT) over the full CoT sequence and the final decision token.
Given a source sentence $x$ and a translation hypothesis $y$, GenRM performs a single long CoT reasoning process that can be abstracted as
$\hat{x}=f_{\mathrm{bt}}(y),\;s(x,\hat{x}),q(y)\in\{0,1\},$ (12)
where $f_{\mathrm{bt}}$ denotes implicit back-translation, s(x,x^)s($\hat{x}$) indicates whether the core information in $x$ is preserved, and $q(y)$ evaluates the intrinsic quality of $y$ in terms of adequacy and fluency. Based on these decisions, GenRM outputs a binary quality reward
$R_{gen}(x,y)=\mathbb{I}\big[s(x,\hat{x})=1\land q(y)=1\big].$ (13)
This formulation yields a reference-free, interpretable quality signal and enforces an explicit conjunction between information preservation and translation quality. Concrete examples of the CoT reasoning are provided in the Appendix H.
4 Experiments and Results
Table 3: Performance Comparison on Sand-Glass Across All Language Pairs (BLEU-$\rho$, Cometkiwi, BT-CERR). All results are averaged by 3 times.
Model Evaluation Metrics on Sand-Glass
Zh $\rightarrow$ En Zh $\rightarrow$ De Zh $\rightarrow$ Es
IB Cometkiwi BT-CERR BLEU-$\rho$ AVG-Tokens IB Cometkiwi BT-CERR BLEU-$\rho$ AVG-Tokens IB Cometkiwi BT-CERR BLEU-$\rho$ AVG-Tokens
Category 1: w/o Syllable Constraint
claude-4.1-opus $\times$ 0.741 0.943 0.313 17.1 $\times$ 0.676 0.882 0.247 24.2 $\times$ 0.668 0.900 0.220 21.3
deepseek-v3 $\times$ 0.748 0.941 0.287 46.3 $\times$ 0.676 0.889 0.222 62.3 $\times$ 0.671 0.884 0.191 99.6
gemini-2.5-pro $\times$ 0.739 0.941 0.307 12.1 (916.9) $\times$ 0.666 0.877 0.238 14.2 (125.8) $\times$ 0.668 0.883 0.212 14.0 (1532.7)
gpt-5 $\times$ 0.746 0.947 0.300 22.4 (412.8) $\times$ 0.681 0.894 0.220 23.2 (605.6) $\times$ 0.671 0.899 0.206 24.0 (676.8)
Category 2: w/ Syllable Constraint
claude-4.1-opus $\checkmark$ 0.691 0.907 0.376 12.2 $\checkmark$ 0.535 0.838 0.293 17.2 $\times$ 0.631 0.849 0.261 13.5
deepseek-v3 $\times$ 0.700 0.924 0.307 12.4 $\times$ 0.557 0.866 0.245 15.8 $\times$ 0.556 0.868 0.225 16.3
gemini-2.5-pro $\times$ 0.698 0.927 0.322 10.7 (1231.3) $\times$ 0.517 0.853 0.266 12.3 (1450.1) $\times$ 0.586 0.827 0.252 10.7 (554.5)
gpt-5 $\times$ 0.705 0.928 0.345 21.7 (2024.9) $\times$ 0.534 0.874 0.283 23.1 (1962.3) $\times$ 0.586 0.833 0.270 21.0 (3105.8)
Category 3: Best of N
claude-4.1-opus $\checkmark$ 0.691 0.903 0.367 109.7 $\checkmark$ 0.611 0.848 0.324 109.2 $\checkmark$ 0.621 0.857 0.304 110.0
deepseek-v3 $\checkmark$ 0.691 0.902 0.332 122.6 $\checkmark$ 0.527 0.863 0.286 160.5 $\checkmark$ 0.522 0.851 0.275 113.6
gemini-2.5-pro $\checkmark$ 0.673 0.904 0.276 133.2 (1089.2) $\checkmark$ 0.517 0.854 0.241 136.0 (972.7) $\checkmark$ 0.554 0.858 0.235 133.8 (611.1)
gpt-5 $\checkmark$ 0.673 0.906 0.296 114.2 (1810.2) $\checkmark$ 0.571 0.861 0.275 111.2 (2229.6) $\checkmark$ 0.529 0.855 0.256 110.8 (2158.8)
Category 4: Translate + Modify Pipeline (2 run)
claude-4.1-opus $\times$ 0.694 0.904 0.327 30.2 $\times$ 0.653 0.868 0.276 42.7 $\times$ 0.643 0.876 0.246 37.6
deepseek-v3 $\times$ 0.703 0.934 0.322 88.1 $\times$ 0.645 0.874 0.239 89.4 $\times$ 0.543 0.873 0.221 102.1
gemini-2.5-pro $\checkmark$ 0.691 0.902 0.348 22.3 (2517.7) $\times$ 0.627 0.830 0.267 27.5 (1043.1) $\checkmark$ 0.602 0.840 0.250 22.4 (2975.2)
gpt-5 $\checkmark$ 0.690 0.917 0.321 42.2 (2666.2) $\checkmark$ 0.514 0.842 0.275 46.6 (2773.0) $\checkmark$ 0.557 0.820 0.234 44.9 (4012.9)
Category 5: HOMURA
Ours (HOMURARubric) $\checkmark$ 0.700 0.914 0.378 9.2 $\checkmark$ 0.591 0.853 0.317 16.3 $\checkmark$ 0.620 0.863 0.260 13.1
Ours (HOMURAReason) $\checkmark$ 0.701 0.925 0.376 10.1 $\checkmark$ 0.603 0.861 0.321 14.7 $\checkmark$ 0.605 0.858 0.309 13.2
Note: IB: In Bounds, $\checkmark$ indicates within bounds, $\times$ indicates out of bounds. AVG-Tokens: Average number of output tokens per sample; for reasoning models (), values in parentheses denote the total tokens consumed (including internal thinking).
4.1 Experimental Setup
We evaluate SOTA LLMs under five strategy categories that span the length-control spectrum:
  • Category 1: Unconstrained LLMs. Frontier models (GPT-5 OpenAI (2025), Claude-4.1-Opus Anthropic (2025), Gemini-2.5-Pro, DeepSeek-V3) with standard decoding. For GPT-5 and Gemini-2.5-Pro, we set low reasoning effort to approximate direct translation. These serve as semantic upper bounds.
  • Category 2: Prompt-based Compression. Category 1 models with brevity-oriented system prompts enforcing syllable-level conciseness.
  • Category 3: Best of N Refinement. Generating $N$ candidates with varying lengths, followed by a greedy selection of the highest-fidelity output that fulfills the syllable budget.
  • Category 4: Pipeline-based Post-editing. Two-stage translation $\rightarrow$ length-constrained rewriting.
  • Category 5: Our Method (HOMURA). RL fine-tuning on Qwen3-8B (Yang et al., 2025) to enforce temporal budgets as hard constraints.
More experimental details are provided in Appendix E.
4.2 Target Syllable Bounds Configuration
Following Section 2.3, we set density-calibrated target bounds $\mathcal{B}_{L}$ to impose comparable compression pressure across languages: En $\in[0.8,0.9]$, De $\in[0.9,1.0]$, and Es $\in[1.0,1.1]$. These intervals are used as hard constraints for HOMURA.
4.2.1 Evaluation Metrics
We report complementary metrics for efficiency, semantic preservation, and linguistic quality:
Translation Density (BLEU-$\rho$ $\uparrow$)
We define BLEU-$\rho$ as semantic fidelity per temporal unit:
$\text{BLEU-}\rho=\frac{\text{BLEU}(x,\hat{x})}{\rho(x,y)},$ (14)
[…]
Semantic Integrity (BT-CERR $\uparrow$)
BT-CERR is a binary indicator that evaluates whether all core events extracted from $x$ are retained in the back-translation $\hat{x}$. (see Appendix D for the definition of core events).
Linguistic Quality (Cometkiwi $\uparrow$)
We use reference-free CometkiwiRei et al. (2022) to assess fluency and adequacy under compression.
4.3 Main Results
4.3.1 Cross-Lingual Performance and Temporal Compliance
Table 3 summarizes performance on Sand-Glass. We report human evaluation in Appendix C and Statistical Significance Testing in Appendix B. Overall, the results highlight the trade-off between temporal compliance and semantic density:
Budget Violation in Frontier Models.
As shown by the IB column, unconstrained frontier model(Cat 1) universally fail to meet temporal budgets ($\times$), particularly in low-density languages like Spanish. While specialized prompting (Cat 2) shifts the distribution toward brevity, it remains unreliable, with models like GPT-5 still exhibiting "verbosity drift" that exceeds target bounds. This confirms that pure instruction-following is insufficient for rigorous syllable-level constraints.
The Search-Fidelity Trade-off.
Search-based (Cat 3) and pipeline (Cat 4) strategies achieve higher In-Bounds rates but at a significant semantic cost. For instance, Cat 4 variants often show a drop in BT-CERR, suggesting that the disjointed post-editing process leads to "hallucinated omissions" of critical source predicates. Moreover, the increased inference latency of Best-of-$N$ makes it less practical in real-time setting.
(a) BLEU vs. Avg. Syllable Ratio
(b) BT-CERR vs. Avg. Syllable Ratio
[…]
Information Density Superiority of HOMURA.
In contrast, HOMURA consistently maintains in bound while achieving the highest BLEU-$\rho$ across all language pairs. This gap demonstrates that HOMURA does not merely truncate text but actively performs semantic packing—delivering a higher "semantic payload" per syllable. Additionally, we scale our method to a 32B model; see Appendix G for details.
Inference efficiency.
HOMURA is markedly more inference-efficient. Unlike Best-of-$N$, which multiplies decoding cost by generating $N$ candidates, and unlike translate-then-modify pipelines that require an extra rewriting pass, HOMURA satisfies the syllable/temporal budget in a single decoding. Consequently, it attains high IB compliance with substantially fewer tokens (AVG-Tokens in Table 3), making it suitable for real-time and high-throughput deployment.
Robustness of Reason-based GenRM.
Within Category 5, HOMURAReason achieves higher BT-CERR than HOMURARubric on two of the three target languages, with the largest gain observed for English. This suggests the reason-based reward better prunes redundancy while preserving the source’s logical structure. We further validate this by replacing the self-trained GenRM in HOMURAReason with an LLM-as-RM baseline (DeepSeek-V3); results are in Appendix J.
Sensitivity of KL Regularization.
While the theoretical objective in Eq. (4) assumes $\beta>0$ for policy stability, our empirical investigations reveal a task-specific trade-off. As shown in our ablation experiments, introducing $\beta>0$ leads to a notable degradation in structural alignment. This suggests that for time-constrained translation, the reference model $P_{0}$ imposes a “verbosity prior” that penalizes the radical lexical pruning and syntactic restructuring necessary to meet the Sand-Glass budget. By setting $\beta\to 0$ in our final implementation, we grant the model maximum compression freedom. The inherent stability of the GRPO algorithm’s group-relative optimization, coupled with the explicit guidance of our task-specific rewards, ensures that this relaxation preserves policy stability and prevents linguistic collapse, allowing HOMURA to reach a more efficient rate-distortion frontier. A detailed theoretical and empirical analysis of this stabilization mechanism is provided in Appendix M.
4.3.2 Analysis of Quality–Compression Trade-off
Figure 4 illustrates the trade-off between translation quality and output length. By plotting semantic metrics against the average syllable ratio $\rho$, we compare how different models perform as the available length budget decreases.
As shown in Figure 4, unconstrained baselines cluster in the high-quality, high-$\rho$ region, representing the model’s performance when length is not a factor. While standard LLMs can be prompted to shorten their outputs, their quality drops sharply as compression becomes more aggressive. This reflects a low rate–distortion efficiency: when forced to reduce the phonetic “rate” (syllables), these models suffer from disproportionately high “distortion” (loss of meaning), as they struggle to identify which information is most essential to preserve.
In contrast, HOMURA variants demonstrate significantly higher rate–distortion efficiency. At any given syllable budget, HOMURA maintains higher semantic fidelity than the baselines; conversely, it can achieve a much shorter length while maintaining the same quality level. These results suggest that our framework does not simply truncate text. Instead, it optimizes the model to operate at a more efficient trade-off point—packing a higher semantic payload into a limited phonetic budget to respect the strict constraints of media localization.
4.3.3 Qualitative Case Studies
We qualitatively compare HOMURA with representative baselines to assess practical behavior under different compression budgets. As shown in Table 10, frontier models (e.g., GPT-5, Claude-4.1-opus) often remain verbose even with brevity prompts, whereas HOMURA achieves tighter budget adherence by packing semantics into denser lexical and syntactic forms. When the budget becomes extreme (down to $\rho\approx 0.25$), HOMURA smoothly shifts from natural prose to highly condensed translations, prioritizing core predicate–argument content while discarding redundant function words. Full case comparisons and a taxonomy of the observed linguistic shifts are deferred to Appendix K.
4.4 Autonomous Compression: Navigating the Compression Limit
The preceding experiments show that HOMURA can reliably translate under a user-specified ratio interval. We further ask whether, without a preset lower bound, the model can autonomously discover the practical compression limit: the point where additional shortening triggers a sharp loss in meaning. To this end, we formulate an autonomous compression task by replacing the hard-interval length reward with a continuous incentive:
$R_{\mathrm{len}}^{\mathrm{(auto)}}(x,y)=\min\left(\cos\left(\theta\cdot\rho(x,y)\right),\;0\right),$ (15)
where $\theta$ scales the incentive gradient. The policy is encouraged to minimize $\rho$ while still being constrained by the quality reward $R_{\mathrm{qual}}$.
Table 4: The Compression Wall. The autonomous model reliably reaches the feasible limit; stricter manual targets miss $\rho$ and produce worse compression.
Model & Configuration $\boldsymbol{\rho}$ BLEU-$\boldsymbol{\rho}$ BT-CERR
Auto HOMURA 0.485 0.404 0.811
Manually Constrained HOMURA
   $R_{\mathrm{target}}\in[0.3,0.4]$ 0.558 0.392 0.819
   $R_{\mathrm{target}}\in[0.4,0.5]$ 0.481 0.384 0.816
   $R_{\mathrm{target}}\in[0.5,0.6]$ 0.559 0.386 0.856
   $R_{\mathrm{target}}\in[0.8,0.9]$ 0.808 0.378 0.914
The Empirical Compression Wall
We observe a stable compression wall when pushing the model to its structural limit. In Table 4, autonomous optimization consistently converges to a narrow operating regime around $\rho\in[0.46,0.49]$, regardless of the specific reward configuration.
[…]
Efficiency at the Frontier
To quantify compression efficiency, we report BLEU-$\rho$. The autonomous model attains the best BLEU-$\rho$ while operating at a genuinely shorter length ($\rho=0.485$ vs. $0.558$), indicating that autonomous optimization finds a more efficient operating point on the quality-compression frontier.
Rate-Distortion Efficiency at the Frontier
Qualitatively, the autonomous model tends to compress by semantic condensation: dropping modifiers, selecting denser lexical items, and simplifying clause structures. In contrast, when forced toward an overly aggressive interval, outputs more often exhibit unstable shortening behaviors, reflecting the tension between further length reduction and preserving core meaning.
Takeaway: The Wall as a Feasibility Boundary
Together, these results suggest an empirical feasibility boundary around $\rho\approx 0.49$ for time-constrained translation in our setting (Zh→En).We hypothesize that this boundary represents the minimal information carrier required to express core predicate-argument content.
5 Conclusion
We address systemic cross-lingual verbosity bias in LLM translation, a key obstacle in time-constrained settings such as subtitles and dubbing. We introduce Sand-Glass, a syllable-budgeted benchmark grounded in the Iso-Information Principle, and propose HOMURA, an RL framework that optimizes the fidelity–feasibility trade-off.
[…]
Limitations
First, while we adopt syllable count as a duration proxy grounded in the Iso-Information Principle, this metric remains a textual approximation of acoustic reality. In real-world dubbing and subtitling, physical duration is heavily influenced by prosodic features and speech tempo, and our current text-only framework does not account for multimodal constraints like lip-synchronization or isochrony. Consequently, generated outputs may still require minor timing adjustments during recording.
Second, our experimental validation is currently restricted to Chinese-to-English, German, and Spanish translation tasks. The universality of the "compression feasibility boundary" observed in our study remains to be verified across linguistically distant pairs or low-resource languages. Future work is needed to determine if such limits are language-specific or intrinsic to the translation task itself.
Acknowledgments
Appendix A Related Work
A.1 Constrained Translation and Subtitling
Media localization imposes strict spatiotemporal constraints beyond semantic fidelity [18]. To address these limitations, early research focused on supervised learning paradigms that integrate specialized markers to guide line length and reading speed [17]. An advancement was introduced by Wu et al. [36] with VideoDubber, which pioneered direct speech-length control by incorporating token-level duration predictions into the decoder stack. However, this approach remains constrained by its dependence on supervised training with non-isochronous datasets and heuristic duration labeling that introduces annotation noise. Building on this, Chadha et al. [3] demonstrated that even with automatically generated length tags derived from phoneme ratios, supervised training on non-isochronous speech-translation data remains a fundamental bottleneck for achieving precise isochronicity. While architectures like CLSC [16] further refine this task using control tokens to manage sequence length, they remain fundamentally dependent on supervised training with task-specific compression corpora. In contrast, our work targets general-purpose Large Language Models (LLMs) via a reinforcement learning framework. This approach explicitly optimizes the trade-off between semantic preservation and temporal compliance without requiring ground-truth compression datasets.
A.2 Hard Constraints and Lyric Translation
The imposition of strict feasibility constraints parallels automatic lyric translation, where models must adhere to syllable counts and rhythm [25]. Building on curriculum RL strategies for discrete constraints [30], HOMURA treats temporal duration as a hard constraint. We employ a KL-regularized objective to balance these rigid length requirements with semantic quality, ensuring the model respects the "sand-glass" temporal budget.
A.3 Verbosity Bias in LLMs
A major challenge for time-constrained translation is the systemic "verbosity bias" in LLMs, often exacerbated by RLHF alignment that conflates length with helpfulness [31, 35, 34]. Our framework counteracts this by introducing a dynamic syllable-ratio reward. Instead of encouraging length drift, we specifically penalize verbosity, aligning the model’s output with the strict density requirements of subtitling and dubbing.
Appendix B Statistical Significance Testing
We perform hypothesis testing to examine whether the improvements brought by HOMURA are statistically significant. For a fair comparison, for each baseline we choose its best-performing configuration that (i) requires a single inference pass and (ii) satisfies the target in-bounds constraint (IB), and compare it with HOMURA under the same constraints. Table 5 reports the resulting p-values.
Table 5: p-values for testing HOMURA against each variant.
Model p-value of HOMURA
Zh $\rightarrow$ En Zh $\rightarrow$ De
IB BLEU-$\rho$ COMETKiwi IB BLEU-$\rho$ COMETKiwi
HOMURARubric $\checkmark$ 6.46e-03 8.20e-03 $\checkmark$ <1e10<\!1e\!-\!10 <1e10<\!1e\!-\!10
HOMURAReason $\checkmark$ - 6.30e-03 $\checkmark$ <1e10<\!1e\!-\!10 <1e10<\!1e\!-\!10
[…]
Appendix C Human Evaluation
We additionally report human evaluation results for the Zh-to-En setting to further validate the effectiveness of our method and the appropriateness of our evaluation metrics.
Table 6: Human evaluation results on Accuracy and Completeness under different settings. Note: IB: In Bounds, $\checkmark$ indicates within bounds, $\times$ indicates out of bounds.
Setting Model IB Accuracy Completeness Avg.
w/o syll. Gemini-2.5-Pro $\times$ 4.95 5.00 4.97
Claude-4.1-Opus $\times$ 4.91 4.98 4.94
GPT-5 $\times$ 4.96 5.00 4.98
DeepSeek-V3 $\times$ 4.97 5.00 4.99
w/ syll. Gemini-2.5-Pro $\times$ 4.44 4.90 4.67
Claude-4.1-Opus $\checkmark$ 4.42 4.93 4.67
GPT-5 $\times$ 4.42 4.87 4.65
DeepSeek-V3 $\times$ 4.42 4.87 4.65
Best-of-NN Gemini-2.5-Pro $\checkmark$ 4.57 4.57 4.57
Claude-4.1-Opus $\checkmark$ 4.51 4.49 4.50
GPT-5 $\checkmark$ 4.71 4.80 4.75
DeepSeek-V3 $\checkmark$ 4.56 4.58 4.57
Ours HOMURARubric $\checkmark$ 4.91 4.83 4.87
HOMURAReason $\checkmark$ 4.85 4.93 4.89
Table 6 shows that (i) without syllable constraints, all frontier LLMs achieve near-ceiling Accuracy/Completeness, serving as an unconstrained upper bound; (ii) enforcing syllable budgets by prompting substantially degrades human-rated quality, especially Accuracy; and (iii) Best-of-$N$ provides limited and unstable gains. In contrast, our HOMURA variants deliver the best overall constrained performance, approaching the unconstrained ceiling, with HOMURARubric slightly favoring Accuracy and HOMURAReason favoring Completeness.
Appendix D The Sand-Glass Benchmark Construction
Sand-Glass is designed for time-constrained translation, where the output must satisfy a strict temporal budget while preserving essential meaning. The construction emphasizes two components: (i) duration-aware segmentation to derive realistic per-segment budgets from natural speech, and (ii) core event extraction to enable a fine-grained semantic retention check under forced compression (facilitating the BT-CERR metric defined in Sec. 4.2.1).
Pipeline Overview.
Starting from real-world video datasets across diverse specialized domains, we extract ASR transcripts, segment them into duration-bounded subtitle units, apply multi-stage quality filtering (including strict length constraints), and finally extract core events (predicate–argument abstractions) to serve as a semantic backbone.
Construction Procedure.
  1. 1.
    Data Acquisition and Segmentation. We collect a large-scale corpus of real-world video transcripts covering five representative domains: Gaming, Film & Television, Travel & Tourism, ACGN (Animation, Comics, Games and Novels), and General Knowledge. Speech is transcribed via ASR and segmented based on pauses and semantic units. Each segment is assigned a temporal budget derived from its actual speech duration.
  2. 2.
    Preprocessing and Packaging. We strip non-speech artifacts (e.g., fillers, music markers) and package segments into sliding windows of ten to preserve contextual coherence. Outputs are enforced into a fixed JSON schema for model interaction.
  3. 3.
    Multi-stage Quality Filtering. We apply a rigorous filtering pipeline to ensure structural and linguistic integrity. Statistical Sanity: To ensure the statistical stability of the $\rho$ metric, we exclude segments with fewer than 10 characters or extreme Characters-Per-Second (CPS) values. This prevents short-sequence outliers from inflating density results and ensures that temporal constraints necessitate meaningful semantic compression. Quality Heuristics: Samples are scored via an ensemble of signals—including perplexity, repetition, and script consistency—retaining borderline cases with lower weights to preserve a natural difficulty curve. Diversity: Following LSH-based de-duplication [14], we perform domain-balanced quota sampling to obtain a high-quality final corpus of 1,000 golden instances. This ensures a representative and manageable distribution of expansion challenges across the five selected domains.
  4. 4.
    Core Event Extraction. Since compression increases the risk of meaning-critical omissions, we extract core events from each source segment as a minimal semantic representation. Concretely, we identify the main predicate and collect its content-bearing arguments, including nouns, proper nouns, and numerals by using using Stanza Chinese POS tagging. [^2] https://stanfordnlp.github.io/stanza/. We then measure semantic consistency by computing contextual embedding similarity between source and candidate predicate–argument realizations, and flag a violation if no candidate event sufficiently matches the source core event. This predicate–argument abstraction serves as the reference for measuring semantic integrity in compressed translations (via BT-CERR).
Appendix E Implementation Details
Our RM training framework is built on Megatron. We use the Qwen3-32B-Chat model as the initialization. The training is conducted with a batch size of 256, using a cosine learning rate scheduler with an initial learning rate of 5e-6. All models are trained on 64 Huawei’s Ascend 910B NPUs.
Our RL training framework is based on the Verl framework. We use the Qwen3-8B-Chat model as the initialization for RL training. During training, we configure a batch size of 16 and perform 16 rollouts per prompt using the GRPO algorithm. The learning rate is initialized at 1e-8, and a cosine scheduler with warm-up is applied toward the final iteration. Sampling is conducted with a temperature of 1.0, and the maximum generation length is limited to 1,024 tokens. The KL penalty coefficient $\beta$ is set to 0, effectively removing the KL constraint relative to the reference policy. The PPO clipping range $\epsilon$ is fixed at 0.2. All models are trained for one epoch using 8 NVIDIA H800 80G GPUs. Other detail training hyperparameters are listed in Table 7.
Our training set consists of 70K instances drawn from the same data sources as the Sand-Glass benchmark, and is constructed following the same procedure.
Table 7: Key hyperparameters used in training HOMURA.
   Hyperparameter    Value
   $\alpha_{1}$    0.4
   $\alpha_{2}$    0.5
   $\lambda_{len}$    0.5
   $\lambda_{qual}$    0.5
   $k$    300
   $\tau_{min}$    0
   $\tau_{max}$    0.8
Appendix F Effectiveness of Dynamic Syllable-ratio Reward
[…]
As shown in Figure 5, the experimental results yield several key observations:
  • Stability and Convergence: The Dynamic reward (blue) demonstrates significantly smoother optimization and faster convergence into the target interval $[0.8,0.9]$. In contrast, the Static reward (red) exhibits violent oscillations, especially in early training phases.
  • Mitigating Discretization Noise: The instability of the static baseline confirms that "hard" rewards for syllable counts are overly sensitive to small variations in short segments. A single-syllable difference in a short sentence can cause a disproportionate jump in the ratio, providing an inconsistent gradient for the RL agent.
  • Optimization Efficiency: By adaptively relaxing the lower bound for shorter sequences via $\gamma(x)$, our method provides a more continuous and achievable reward landscape. This prevents the model from being penalized for unavoidable linguistic granularity, leading to more efficient learning of the desired compression behavior.
These findings justify the use of length-aware dynamic bounds in Homura to ensure stable policy refinement across diverse utterance lengths.
Appendix G Scaling to Larger Model.
[…]
Scale Model IB Cometkiwi BT-CERR BLEU-$\rho$
8B HOMURARubric $\checkmark$ 0.700 0.914 0.378
HOMURAReason $\checkmark$ 0.701 0.925 0.376
32B HOMURARubric $\checkmark$ 0.740 0.917 0.388
HOMURAReason $\checkmark$ 0.740 0.933 0.385
We analyze the effect of model scaling by comparing 8B and 32B variants of HOMURA under identical training settings, with results reported in Table 8. Scaling consistently improves translation quality and semantic preservation, as shown by higher BLEU-$\rho$, Cometkiwi and BT-CERR scores. Moreover, the relative advantage of HOMURAReason over HOMURARubric is preserved at larger scale, suggesting that the proposed reward design scales robustly with model capacity.
Appendix H Examples of Structured CoT Reasoning for GenRM
This appendix provides illustrative examples of the structured CoT reasoning produced by GenRM.
Structured CoT Example #1 Source sentence (x):
我曾经在这看到过几艘陌生的船。
Translation hypothesis (y):
I have seen a few strange ships here before.
COT process: 1. Back-translation: The translation “I have seen a few strange ships here before.” is back-translated into Chinese. The back-translation result is “我以前曾在这里见过几艘奇怪的船。” (I have previously seen several strange ships here). Alternatively, it can also be “我曾经在这里看到过几艘陌生的船。” (I once saw several unfamiliar ships here). 2. Semantic consistency judgment: Compare the source sentence “我曾经在这看到过几艘陌生的船。” (I once saw several unfamiliar ships here) with the back-translated content “我以前曾在这里见过几艘奇怪的船。” (I have previously seen several strange ships here). The two are completely consistent in core semantics, both expressing the fact of “过去在某地看到过不熟悉的船只” (having seen unfamiliar ships at a certain place in the past). “曾经” (once / before) corresponds to “before,” “在这” (here) corresponds to “here,” “看到过” (have seen) corresponds to “have seen,” “几艘” (a few) corresponds to “a few,” “陌生的” (unfamiliar / strange) corresponds to “strange,” and “船” (ships) corresponds to “ships.” The semantic consistency is very high. 3. Translation quality judgment: The translation “I have seen a few strange ships here before.” is grammatically correct, using the present perfect tense to express a past experience, which conforms to English usage. The expression is fluent and natural, with accurate word choice. Therefore, the translation quality is high. Based on the above, the semantics are consistent and the translation quality is high, resulting in a final score of 1. Score:
1
Structured CoT Example #2 Source sentence (x):
所以天命人只有超度了金池长老的冤魂
Translation hypothesis (y):
So the Mandate Person can only rescue the lingering soul of Jinchi Elder
COT process: 1. Back-translation: The translation “So the Mandate Person can only rescue the lingering soul of Jinchi Elder” is back-translated into Chinese as “所以天命人只能解救金池长老逗留的灵魂。” (So the Mandate Person can only rescue the lingering soul of Elder Jinchi). 2. Semantic consistency judgment: The key terms in the source text are “超度” (perform salvation rites / help a soul attain deliverance) and “冤魂” (wronged or vengeful spirit). “超度” is a term with strong religious and cultural connotations, referring to helping the deceased soul attain liberation and enter reincarnation or the afterlife through rituals. However, the word “rescue” in the translation is overly general and fails to convey the deeper religious meaning of “超度”. “冤魂” refers to a ghost that died unjustly and carries strong resentment, whereas “lingering soul” (逗留的灵魂) only expresses the idea of lingering or wandering, losing the core semantic component of “冤” (wronged / resentful). Therefore, the back-translated content is semantically inconsistent with the original text. 3. Translation quality judgment: Due to the semantic inconsistency, this step is skipped. Based on the above analysis, the semantics are inconsistent, and the final score is 0. Score:
0
Appendix I Prompt for Linguistic well-formedness Reward.
This appendix provides the training prompt used for the linguistic well-formedness reward.
Training Prompt You are an expert in {target_lang} translation. Given context <context> for background reference. Here is a text <text> from the context and its corresponding {target_lang} translation <translation>. You need to determine whether <translation> is a qualified {target_lang} translation. Requirements: Consider only the following criteria; if any are not met, the score is 0: 1. Grammaticality: No obvious grammatical or syntactic errors, consistency errors (e.g., subject–verb agreement, tense, case, gender/number), collocation errors, or linguistic defects caused by punctuation or spelling. 2. Readability & Coherence: The sentence structure is complete, logically coherent, and naturally connected, making it smooth and easy to understand; there should be no obvious fragmentation, garbled text, repetitive stacking, or malformed sentences. Input: <context>: {context} <text>: {text} <translation>: {translation} Output Requirements: If <translation> satisfies both criteria above, return <<1>>. Otherwise, return <<0>>. Please output <<0>> or <<1>> directly; do not output any explanation. Output:
Appendix J Ablation of GenRM Design
This appendix presents an ablation study on the design of the reward model within the GenRM framework. In particular, we compare our self-trained GenRM with an external LLM-as-RM. In this ablation, the external reward model fully replaces GenRM, while all other components of the compression-oriented reinforcement learning pipeline are kept identical.
External LLM-as-RM Setup
For the external reward model baseline, we adopt an LLM-as-RM setup using DeepSeek-V3. Given a source sentence and its translation, the model is prompted to perform a quality assessment. Specifically, it evaluates whether the translation (i) preserves the core semantic content of the source sentence under the target syllable constraint, and (ii) remains fluent and grammatical. The model outputs a binary acceptability judgment, which is directly used as the reward signal during policy optimization. The prompt for external LLM-as-RM is below:
Prompt for GenRM You are a translation quality reward model (GenRM). Given contextual information, a source sentence, and its translation, you will perform step-by-step reasoning to assess translation quality. Your evaluation process must strictly follow the sequence below. Step 1: Back-Translation Translate the given translation back into the source language. Step 2: Semantic Consistency Assessment Compare the semantic consistency between the original source text and the back-translated text. If the meanings are consistent or highly consistent, proceed to Step 3. If they are inconsistent, skip Step 3 and assign a final score of 0. Step 3: Translation Quality Assessment Evaluate whether the translation is of high quality based on factors such as fluency, grammatical correctness, and cultural appropriateness. Your final output must include a clear chain-of-thought reasoning process and a reward score: If semantic consistency is satisfied and the translation quality is high, set score = 1. Otherwise, set score = 0. Output Format The output must strictly follow the format below. Do not add or omit any fields: {
"COT": "<COT reasoning>",
"score": 0 or 1
Input Format Context: {context} Source text: {current_text} Translation: {translated_text}
Results and Analysis
Table 9 compares compression translation models optimized with different reward model designs. The model trained with the self-trained GenRM consistently outperforms its counterpart using an external LLM-as-RM baseline, demonstrating that a reward model trained specifically for compression-oriented translation provides more effective optimization guidance than a generic, prompt-based evaluator.
Table 9: Effect of GenRM choice on compressed translation under HOMURAReason (Average across all language pairs).
Model Variant Cometkiwi BT-CERR BLEU-$\rho$
HOMURAReason 0.636 0.881 0.335
HOMURAReason (w/ External RM) 0.628 0.879 0.330
Appendix K Case Study
This appendix provides a fine-grained analysis of the linguistic strategies employed by Homura to navigate the quality-compression frontier. Unlike standard LLMs that often rely on simple truncation, our model learns a hierarchy of compression operations.
Observation 1: Taxonomic Shifts in Semantic Packing.
Analysis of Table 10 reveals three distinct tiers of semantic packing performed by Homura as constraints tighten:
  • Lexical Consolidation: Mapping verbose multi-word phrases to high-density synonyms (e.g., the 7-syllable “relationship between them” $\rightarrow$ 1-syllable “bond”).
  • Aspectual Simplification: Reducing periphrastic verbal constructions into synthetic forms (e.g., “is becoming closer” $\rightarrow$ “grows closer”), which preserves the temporal aspect while reducing syllable count.
  • Syntactic Pruning: Selectively removing low-surprisal functional tokens (articles, auxiliaries) while anchoring the sentence around core predicate-argument structures.
Observation 2: Behavioral Transition at the Compression Wall.
The trade-off spectrum in Part 2 of Table 10 visualizes the transition from naturalistic translation to “telegraphic speech.” While fluency is maintained down to $\rho\approx 0.50$, pushing toward the empirical limit ($\rho\approx 0.25$) forces the model to prioritize propositional content over morphosyntactic correctness. This behavior confirms that HOMURA does not randomly drop tokens but strategically re-allocates the syllable budget to the most informative constituents (e.g., “Bond grows tighter”), ensuring minimal meaning loss even at the threshold of reward collapse.
Table 10: Qualitative Case Study on $\text{Zh}\rightarrow\text{En}$: Comparison between frontier LLMs and Homura across varying compression intensities. Source: 两人之间的关系越来越亲密 (12 syllables).
Category Model Output Text Syl. $\boldsymbol{\rho}$ Quality Observation
Part 1: Comparison with Baselines (Target $\rho\in[0.8,0.9]$ for Prompted Models)
Unconstrained Claude-4.1-Opus The relationship between the two is becoming increasingly intimate. 20 1.67 Naturally verbose; high fidelity.
Gemini-2.5-Pro The relationship between the two grew increasingly intimate. 17 1.42 Fluent but ignores brevity.
GPT-5 The relationship between the two is growing increasingly close. 17 1.42 Optimal quality; no length control.
Prompt-based Compression Claude-4.1-Opus Their relationship grows increasingly intimate. 13 1.08 Fails to reach target interval.
Gemini-2.5-Pro Their relationship is growing more intimate. 12 1.00 Fails to reach target interval.
GPT-5 The two are becoming more and more intimate. 12 1.00 Fails to reach target interval.
Part 2: Controllable Compression via Homura ($\rho$ Trade-off Analysis)
Homura $\rho\approx 1.25$ The relationship between the two is becoming closer. 15 1.25 Full fidelity and naturalness.
$\rho\approx 0.83$ The bond between them is growing closer. 10 0.83 Precise adherence via condensation.
$\rho\approx 0.58$ Bond between them grows closer. 7 0.58 Minor grammatical omission.
$\rho\approx 0.50$ Their bond grows closer. 6 0.50 Extreme brevity; core meaning only.
$\rho\approx 0.25$ Bond grows tighter. 3 0.25 Compression Limit; loss of fluency.
Appendix L Detailed Breakdown of Expansion Metrics
Figure 6: Statistical distribution of translation expansion metrics across DE, EN, and ES. The boxplots illustrate the variations in Forward, Back, and Roundtrip expansion ratios, with diamonds indicating means.
To provide a comprehensive view of the verbosity bias identified in Section 2, we present a fine-grained statistical breakdown of expansion metrics across various frontier LLMs and language pairs. As summarized in Table 11, the Roundtrip Expansion Ratio ($R_{rtp}$) consistently exceeds the unity baseline across all tested models and languages, with a high percentage of segments ($60\%$) exhibiting systemic inflation.
The distribution shifts visualized in Figure 6 further confirm that this bias is not merely a result of specific difficult cases but represents a structural characteristic of current LLM-based translation.
Table 11: Detailed Breakdown of Expansion Metrics across Models and Language Pairs. $R_{fwd}$, $R_{bwd}$, and $R_{rtp}$ denote Forward, Backward (Back-translation), and Roundtrip Expansion Ratios, respectively. Values are presented as Mean and Standard Deviation.
Language Model $R_{fwd}$ (Mean $\pm$ Std) $R_{bwd}$ (Mean $\pm$ Std) $R_{rtp}$ (Mean $\pm$ Std) $R_{rtp}>1$ (%)
De claude-4.1-opus $1.58\pm 0.96$ $0.95\pm 1.54$ $1.39\pm 0.99$ 66.2
deepseek-v3 $1.67\pm 1.37$ $0.91\pm 0.29$ $1.54\pm 1.89$ 63.8
gemini-2.5-pro $1.51\pm 0.88$ $0.96\pm 0.44$ $1.35\pm 0.84$ 59.0
gpt-5 $1.63\pm 0.96$ $0.90\pm 0.19$ $1.43\pm 0.87$ 71.8
En claude-4.1-opus $1.36\pm 1.13$ $1.07\pm 0.55$ $1.41\pm 1.32$ 64.8
deepseek-v3 $1.41\pm 1.15$ $1.05\pm 0.48$ $1.47\pm 1.71$ 59.1
gemini-2.5-pro $1.29\pm 0.81$ $1.07\pm 0.25$ $1.33\pm 0.80$ 61.0
gpt-5 $1.33\pm 0.81$ $1.07\pm 0.23$ $1.36\pm 0.80$ 65.9
Es claude-4.1-opus $1.81\pm 1.05$ $0.90\pm 2.48$ $1.41\pm 1.19$ 66.9
deepseek-v3 $1.99\pm 1.83$ $1.01\pm 1.12$ $1.99\pm 2.90$ 64.0
gemini-2.5-pro $1.72\pm 0.99$ $0.83\pm 0.36$ $1.35\pm 0.87$ 60.4
gpt-5 $1.86\pm 1.09$ $0.78\pm 0.18$ $1.40\pm 0.83$ 68.5
Appendix M Theoretical and Empirical Analysis of Regularization Strength
In Eq.(4), we defined the KL-regularized objective where $\beta$ controls the trade-off between control strength and faithfulness to the base model $P_{0}$. In our experiments, we observed that setting $\beta\to 0$ leads to superior syllable-ratio compliance and linguistic quality. This appendix provides a rigorous justification for why the stability typically guaranteed by explicit KL regularization is preserved in HOMURA through the GRPO framework, even in the control-only regime ($\beta=0$).
M.1 Theoretical Justification: From Token-level to Sequence-level Regularization
Standard RLHF objectives incorporate a token-level KL penalty to anchor the policy to a reference model. As shown in Eq.(4), this term imposes a local constraint on every transition, forcing $\pi_{\theta}$ to minimize divergence from the prior of $P_{0}$ at each step. For time-constrained translation (Sand-Glass), satisfying the control condition $c$ requires radical lexical pruning and syntactic restructuring. Under these conditions, the KL term acts as high-dimensional friction, making structural deviations mathematically expensive and hindering the model from reaching the optimal rate-distortion frontier.
By setting $\beta=0$, we prioritize control strength over faithfulness, granting the model maximum compression freedom. Numerical stability is maintained through the synergy of two stabilizing forces:
  • Algorithmic Stability via Centered Clipped Objective: Stability is maintained through a surrogate loss function $\mathcal{L}_{GRPO}$ that integrates group-relative normalization with a clipping mechanism. The policy update is derived from:
    $\mathcal{L}_{GRPO}(\theta)=\frac{1}{G}\sum_{i=1}^{G}\sum_{t=1}^{T}\min\left(r_{i,t}(\theta)\hat{A}_{i},\text{clip}(r_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i}\right)$ (16)
    where $r_{i,t}(\theta)$ is the token-level probability ratio, and the advantage $\hat{A}_{i}$ is normalized within the group:
    $\hat{A}_{i}=\frac{R(x,c,y_{i})-\mu_{group}}{\varsigma_{group}},\quad\mu_{group}=\frac{1}{G}\sum_{j=1}^{G}R(x,c,y_{j})$ (17)
    This formulation ensures stability via (1) Relative Anchoring, where the zero-sum nature ($\sum\hat{A}_{i}=0$) prevents global policy drift by centering updates around the group’s moving average; and (2) Trust Region Enforcement, where clipping mitigates policy staleness [39]. Together, they preserve the validity of the first-order approximation for sequence-level rewards without requiring an explicit KL penalty.
  • Implicit Structural Regularization: The syllable-ratio reward $R_{\rho}$ itself functions as a structural regulator. By penalizing linguistic redundancy, $R_{\rho}$ induces a form of structural sparsity, forcing the model to operate within a lower-dimensional manifold of the total policy space. This task-specific constraint effectively prunes suboptimal trajectories that are semantically valid but temporally excessive, providing an alternative stabilization force that aligns the optimization path directly with the Sand-Glass objective.
M.2 Empirical Analysis
To validate these theoretical insights, we analyze the training process across different KL coefficients (Figure 7).
Figure 7: Training dynamics across $\beta\in\{0,0.01,0.05\}$. From left to right: (1) KL Loss, (2) Entropy Loss, (3) Policy Gradient Loss, (4) Gradient Norm, and (5) Syllable Ratio. The $\beta=0$ variant (light blue) achieves the best balance between exploration and stability.
Dynamics and Gradient Vitality.
As illustrated in Figure 7, the $\beta=0$ regime (light blue) demonstrates superior optimization vitality. By removing the faithfulness constraint, the model maintains a healthy and stable Gradient Norm ($\sim$4.0) and robust Policy Gradient Loss, confirming that reward signals are effectively translated into policy updates without being neutralized by optimization friction.
Notably, while $\beta=0$ maintains a high and purposeful Policy Entropy (signifying broad compression-driven exploration), the $\beta=0.05$ variant (purple) exhibits an abnormal surge in entropy mid-training. We interpret this as optimization paralysis: the excessive KL penalty forces the probability distribution to flatten as the model becomes indecisive between length compliance and faithfulness to the verbose prior. This conflict leads to a stagnating Gradient Norm in $\beta>0$ variants, where the model settles for conservative, localized edits once it hits the KL-constrained boundary.
Quantitative Performance: Non-linear Trade-offs.
Table 12 highlights a non-linear relationship between KL regularization and performance. While all configurations successfully satisfy the syllable budget, their linguistic strategies differ:
Table 12: Impact of $\beta$ on final performance. Reaching the budget through $\beta=0$ preserves significantly more linguistic structure (CERR) by favoring structural adaptation over conservative edits. IB denotes the inclusion of the Information Bottleneck module.
Configuration IB COMET BT-CERR BLEU-$\rho$
$\beta=0$ $\checkmark$ 0.700 0.914 0.378
$\beta=0.01$ $\checkmark$ 0.708 0.874 0.332
$\beta=0.05$ $\checkmark$ 0.711 0.879 0.359
  • The Superiority of $\beta=0$: The $\beta=0$ variant achieves the highest CERR and BLEU-$\rho$, significantly outperforming all regularized settings. This confirms that total liberation from the token-level prior is essential for structural adaptation—the model’s ability to fundamentally rewrite sentences for conciseness rather than merely deleting tokens.
  • The Conservative Trap of $\beta>0$: Interestingly, while $\beta=0.05$ shows a slight recovery in COMET compared to $\beta=0.01$, its CERR remains substantially lower than the $\beta=0$ baseline. This suggests that higher $\beta$ values force the model into a conservative pruning regime: it clings to the lexical surface of the reference to satisfy the KL penalty (thereby maintaining COMET levels), but at the cost of failing to find the optimal structural alignment (low CERR and BLEU-$\rho$), eventually damaging the natural flow of the synchronized output.
The data confirms that the lack of optimization momentum (decaying Gradient Norm in Figure 7) for $\beta>0$ prevents the model from escaping the reference model’s verbosity trap. By setting $\beta=0$, the policy can effectively navigate the complex rate-distortion frontier required for high-precision synchronization.
Appendix N Declaration of AI Assistance
To enhance development efficiency, GitHub Copilot was utilized as a coding assistant in this study. Specifically, it facilitated the implementation of data preprocessing modules, the execution of statistical analyses, and the rendering of complex experimental figures. The core algorithmic design, reinforcement learning frameworks, and the interpretation of results remain the original work of the authors, who maintain full responsibility for the code’s accuracy.